OpenStudy (anonymous):

Jack usually mows his lawn in 4 hours. Marilyn can mow the same yard in 3 hours. How much time would it take for them to mow together. Please help!

OpenStudy (anonymous):

in one hour jack mows = 1/4 lawn in one hour marilyn mows = 1/3 lawn both in 1 hour = 1/4 +1/3 = 7/12 lawn 7/12 lawn in 1 hour , so 1 lawn in 12/7 hours

OpenStudy (anonymous):

I need a good answer please

OpenStudy (anonymous):

in 1 hour Jack mows 1/4 of lawn In 1 hour Marilyn can mow 1/3 Together in 1 hour they can mow 1/4 + 1/3 of lawn = 7/12 of lawn so to mow the full lawn together they will take 12/7 hours

OpenStudy (anonymous):

yes what i sed....????

OpenStudy (anonymous):

u r correct Randi !!

OpenStudy (anonymous):

H@H@ thank you lolz

OpenStudy (anonymous):

How come u r able to solve such questions but are putting your entire algebra homework on the site for solution ???

OpenStudy (anonymous):

because I am having a hard time figuring them out, I have been studying for them since this morning and my brain is UUGGHH of equations like these, these answers arn't equations lolz

OpenStudy (anonymous):

and my homework has 100 questions on it and its krazie!!!! lolz

OpenStudy (anonymous):

I suggest that you do algebra on your own and ask ppl to check yr answers n confirm that they r ok Otherwise your grounding in algebra will suffer and it will affect other areas of sciences too...

OpenStudy (anonymous):

alright :D